V = xy+ λ(2,000-20x-10y)
where λ is the Lagrange multiplier.
Now, the first-order conditions for constrained output maximisation are
how i slove it
1
Expert's answer
2020-02-03T05:13:29-0500

"\nabla_{x,y,\lambda} =(\frac{\partial V}{\partial x},\frac{\partial V}{\partial y},
\frac{\partial V}{\partial x}) =0"


Take partial derivatives of "V(x,y,\lambda)= xy+\lambda(2,000-20x-10y)"

"\frac{\partial V(x,t,\lambda)}{\partial x} = y-20\lambda"

"\frac{\partial V(x,t,\lambda)}{\partial y} = x-10\lambda"

"\frac{\partial V(x,t,\lambda)}{\partial \lambda} = 2,000-20x-10y"

Finally, we get the system of equations

"\begin{cases} y-20\lambda = 0 \\ x-10 \lambda=0\\ 2,000-20x-10y=0\end{cases}"

Solve this system:

"\begin{cases} y=20\lambda \\ x=10 \lambda\\ 2,000-200\lambda-200\lambda=0\end{cases}"


"\begin{cases} y=20\lambda \\ x=10 \lambda\\ 400\lambda= 2,000\end{cases}"


"\begin{cases} \lambda= 5 \\ x=50 \\y=100\end{cases}"